One plane took off from the San Fransisco , traveling east. At the same time another plane took off from the same airport, traveling west. The plane traveling east was going 120 miles per hour faster than the plane traveling west. After three hours, the planes were 2,400 miles apart. How fast was each plane going?
Answer by mananth(16949)

traveling east. speed= 120 miles per hour faster than the plane traveling west.
After three hours, the planes were 2,400 miles apart. How fast was each
let speed of west bound plane be x mph
east bound plane speed will be x-120 mph
..
distance = speed * time
time = 3hours
distance traveled by one plane = 3x
distance by other plane = (x-120)*3
..
3x+3(x-120)=2400
3x+3x-360=2400
6x=2760
x=460 speed of west bound plane.
East bound plane speed =460-120= 340 mph
CHECK effective speed = 340+460 = 800 mph
distance = 2400 miles
time = distance / speed = 2400/800 = 3 hours